The standard deviation of the breaking strengths of certain cables produced by a company is given as 240 lbs. After a change was introduced in the process of manufacture of these cables, the breaking strengths of a sample of 8 cables showed a standard deviation of 300 lbs. investigate the significance of the apparent increase in variability, using a significance level of (a) 0.05 and (b)0.01.
1
Expert's answer
2020-05-12T16:22:33-0400

"\sigma_0=240\\
N=8\\
s=300\\
\text{We assume that the breaking strengths has normal distribution}.\\
a)\alpha=0.05\\
H_0:\sigma^2=\sigma_0^2=240^2, H_1: \sigma^2\neq\sigma_0^2=240^2\\
\text{We will use the following random variable:}\\
\chi^2=\frac{(N-1)s^2}{\sigma_0^2}\text{ where } S^2\text{ is a random value}\\
\text{of corrected variance}.\\
\text{This random variable has }\chi^2\text{-distribution with } \\
k=N-1=8-1=7 \text{ degree of freedom}.\\
\text{Observed value:}\\
\chi^2\approx 10.9375.\\
\text{Critical values:}\\
\chi^2_{critical_{left}}=\chi^2_{critical}(1-\alpha/2;k)\approx 1.6899.\\
\chi^2_{critical_{right}}=\chi^2_{critical}(\alpha/2;k)\approx 16.013.\\
\text{Critical region: }(-\infty;1.6899)\cup (16.013;\infty)\\
\text{Our observed value does not fall into the critical region}.\\
\text{So we accept } H_0.\\
\text{Variability did not change after a change}\\
\text{in the process of manufacture}.\\
b)\alpha=0.01\\
\text{Critical values:}\\
\chi^2_{critical_{left}}\approx 0.9893\\
\chi^2_{critical_{right}}\approx 20.278\\
\text{Critical region: }(-\infty;0.9893)\cup (20.278;\infty)\\
\text{Our critical value does not fall into the critical region.}\\
\text{So we accept } H_0.\\
\text{Variability did not change after a change}\\
\text{in the process of manufacture}."